# Large Language Models for Cyber Security: A Systematic Literature Review

### v2

@@ 4 The types and features of datasets used
3. RQ1: What LLMs have been employed to support security tasks?
3.1. Architecture of LLMs in Security
Pre-trained Language Models (PLMs) have exhibited impressive capabilities across various NLP tasks (Shanahan, 2023; Kojima et al., 2023; Minaee et al., 2024; Wei et al., 2023a; Zhao et al., 2023). Researchers have noted substantial improvements in their performance as model size increases, with surpassing certain parameter thresholds leading to significant performance gains  (Shanahan, 2023; Hoffmann et al., 2022). The term ”Large Language Model” (LLM) distinguishes language models based on the size of their parameters, specifically referring to large-sized PLMs (Minaee et al., 2024; Zhao et al., 2023).However, there is no formal consensus in the academic community regarding the minimum parameter size for LLMs, as model capacity is intricately linked to training data size and overall computational resources (Kaplan et al., 2020). In this study, we adopt to the LLM categorization framework introduced by Panet et al. (Pan et al., 2024), which classifies the predominant LLMs explored in our research into three architectural categories: encoder-only, encoder-decoder, and decoder-only. We also considered whether the related models are open-source. Open-source models offer higher flexibility and can acquire new knowledge through fine-tuning on specific tasks based on pre-trained models, while closed-source models can be directly called via APIs, reducing hardware expenses. This taxonomy and relevant models are shown in Table 4.
Model Release Time Open Source
Encoder-Only BERT (8) 2018.10 Yes
RoBERTa (12) 2019.07 Yes
DistilBERT (3) 2019.10 Yes
CodeBERT (8) 2020.02 Yes
DeBERTa (1) 2020.06 Yes
GraphCodeBERT (1) 2020.09 Yes
CharBERT (1) 2020.11 Yes
SecureBERT (1) 2022.04 Yes
Encoder-Decoder T5 (4) 2019.10 Yes
BART (1) 2019.10 Yes
PLBART (3) 2021.03 Yes
CodeT5 (5) 2021.09 Yes
UniXcoder (1) 2022.03 Yes
Flan-T5 (1) 2022.10 Yes
Decoder-Only GPT-2 (9) 2019.02 Yes
GPT-3 (4) 2020.04 Yes
GPT-Neo (1) 2021.03 Yes
CodeX (9) 2021.07 No
CodeGen (5) 2022.03 Yes
InCoder (1) 2022.04 Yes
PaLM (3) 2022.04 No
Jurassic-1 (1) 2022.04 No
GPT-3.5 (52) 2022.11 No
LLaMa (4) 2023.02 Yes
GPT-4 (38) 2023.03 No
Bard (8) 2023.03 No
Claude (3) 2023.03 No
StarCoder (3) 2023.05 Yes
Falcon (2) 2023.06 Yes
CodeLLaMa (4) 2023.08 Yes
Encoder-only LLMs. Encoder-only models, as their name implies, comprise solely an encoder network. Initially designed for language understanding tasks like text classification, these models, such as BERT and its variants (Aghaei et al., 2023b; He et al., 2021; Devlin et al., 2019; Liu et al., 2019; Ma et al., 2020; Feng et al., 2020; Guo et al., 2021; Sanh et al., 2020), aim to predict a class label for input text (Devlin et al., 2019). For instance, BERT, which adopts the encoder architecture of the Transformer model, is mentioned in 35 papers included in this study. During training, its unique bidirectional attention mechanism enables it to simultaneously consider both the left and right context of each word. In the realm of security, various prominent models such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), CharBERT (Ma et al., 2020), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2020) have gained widespread usage. Models like SecureBERT (Aghaei et al., 2023b), tailored for cybersecurity applications, have emerged, innovating upon BERT by incorporating program structure, introducing new pre-training tasks, or integrating different modalities. This enhances the model’s feature representation capabilities and boosts its performance in security-related tasks. An example is RoBERTa (Liu et al., 2019), which enhances BERT’s robustness through various model design adjustments and training techniques. These include altering key hyperparameters, eliminating the next-sentence pre-training objective, and utilizing substantially larger mini-batches and learning rates during training. CodeBERT (Feng et al., 2020) is a bimodal extension of BERT that utilizes both natural language and source code as its input. It employs a replaced token detection task to bolster its understanding of programming languages, in order to tackle code generation and vulnerability detection tasks. Encoder-only LLMs have demonstrated their effectiveness in tasks that require understanding of text or code snippets, such as threat analysis (Aghaei et al., 2023a), intrusion detection (Aghaei and Al-Shaer, 2023), vulnerability detection (Thapa et al., 2022), and bug repair (Zhang et al., 2023c) among other security tasks (Manocchio et al., 2024; Fayyazi and Yang, 2023; Aghaei et al., 2023b; Chen et al., 2023b; Xia and Zhang, 2023b; Du and Yu, 2023).
Encoder-decoder LLMs. The Transformer model, based on the encoder-decoder architecture (Vaswani et al., 2023), consists of two sets of Transformer blocks: the encoder and decoder. Stacked multi-head self-attention layers are used by the encoder to encode the input sequence, generating latent representations. In contrast, the decoder performs cross-attention on these representations and sequentially produces the target sequence. Models like BART (Lewis et al., 2019), T5 (Raffel et al., 2023), and CodeT5 (Wang et al., 2021) exemplify this architecture. CodeT5 (Wang et al., 2021) and PLBART (Ahmad et al., 2021) have built upon the foundation of their original models by introducing bimodal inputs of programming language and text, demonstrating effective code comprehension capabilities. Raffle et al. (Raffel et al., 2023) show that almost all NLP tasks can be framed as a sequence-to-sequence generation task in their work. In the domain of security, the flexible training strategy of the encoder architecture, along with its generality for multiple tasks, has been applied to solve issues in web fuzzing, program fuzzing, and reverse engineering (Liu et al., 2020a; Liang et al., 2024; Deng et al., 2023c; Armengol-Estapé et al., 2024).
Decoder-only LLMs. Unlike the encoder-decoder architecture, which involves the encoder processing input text and the decoder generating output text by predicting subsequent tokens from an initial state, decoder-only LLMs rely solely on the decoder module to produce the target output text (Radford et al., 2018). This training paradigm prioritizes sequential prediction and heavily depends on the model’s capacity to comprehend and forecast the structure, syntax, and context of language. GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023b) belong to the GPT series of models, among which GPT-3.5 and GPT-4 are the models most frequently used to address various security issues in this study, covering almost all security application directions(Ferrag et al., 2024; Fayyazi and Yang, 2023; Scanlon et al., 2023; Deng et al., 2023a). More specialized versions like Codex (Chen et al., 2021) and others have been fine-tuned for specific code related tasks. Open-source models like GPT-Neo (Black et al., 2021), LLaMa (Touvron et al., 2023), and Falcon (Penedo et al., 2023) also follow this architecture. Additionally, code generation LLMs such as CodeGen (Nijkamp et al., 2023), InCoder (Fried et al., 2023), StarCoder (Li et al., 2023a), and CodeLLaMa (Rozière et al., 2024) have been widely used to address bug detection and repair, as well as vulnerability repair issues (Wu et al., 2023a; Xia and Zhang, 2023b; Yang et al., 2023c). Decoder-only LLMs have gained prominence in recent years, especially in 2022 and 2023, witnessing a surge in development and commercial adoption by leading Internet companies. For instance, Google introduced Bard (Google, 2023), while Meta unveiled LLaMa (Touvron et al., 2023). Unlike GPT-4 and its derivative application, ChatGPT, which quickly found integration into various security tasks, these newer models have yet to see widespread adoption in the security domain. Their potential remains largely untapped, presenting opportunities for further evaluation and utilization in specific challenges and tasks. The ongoing progress of these models underscores the active exploration and innovation within decoder-only architectures, highlighting their value as tools in security research due to their ability to perform downstream tasks with minimal fine-tuning or additional heads.
3.2. Trend Analysis
Figure 4. Trend of different model architectures.
Illustrated in Figure 4, from 2020 to 2024, there have been significant shifts in the preference and utilization of LLM architectures across security tasks. The selection among decoder-only, encoder-decoder, and encoder-only structures has influenced diverse research directions and solutions in the security field. This examination delves into the trends regarding the adoption of these architectures over time, reflecting the evolving landscape of LLM applications for security tasks.
Evolution of LLM architectures in 2022. In the years 2020 and 2021, only 5 research papers applied LLMs to security tasks. In 2020, one paper each utilized encoder-only and encoder-decoder LLMs to address security issues. However, the situation reversed in 2021, with only decoder-only LLMs being used. We presume this is due to researchers in 2020 and 2021 focusing more on LLMs’ performance in text-related tasks as language models and on updates and innovations in LLM technology, rather than on their applications in other fields (Kaplan et al., 2020; Brown et al., 2020). In 2022, there was a significant shift, as the number of papers employing LLMs for security tasks rose to 12, surpassing the combined total of related papers from the previous two years. Additionally, 2022 witnessed a notable increase in diversity, with a broader range of LLM architectures being represented. Out of the total 12 papers, 5 utilized encoder-only LLMs, constituting 41.7% of the research. Encoder-decoder LLMs were featured in 3 papers, accounting for 25%. Meanwhile, decoder-only LLMs garnered 33% of the research interest with 4 papers. This varied distribution indicates that researchers are actively exploring and utilizing different architectures to address various security needs and challenges. The near equal interest in different architectures underscores the richness of the field, suggesting that no single approach has emerged as the definitive choice.
Dominance of the decoder-only architecture in 2023 and 2024. 2023 and 2024 signaled a strong shift towards decoder-only LLMs, decoder-only LLMs have begun to take a leading role in the application of LLMs to solve security issues. This is closely related to the powerful text comprehension and reasoning capabilities (Ouyang et al., 2022; Wei et al., 2023b) demonstrated by chatbots like ChatGPT, which are based on a decoder-only architecture. These models require minimal fine-tuning and are capable not only of generating syntactically correct code snippets but also functionally relevant ones (Laskar et al., 2023a; Sadik et al., 2023). Their proficiency in understanding the context of the code quickly made them the preferred choice. These papers focus on decoder-only LLMs, accounting for 69.67% of total research in 2023. In contrast, encoder-decoder LLMs were the subject of 14 papers, contributing 11.48% of the papers, while encoder-only LLMs had 23 papers, making up 18.85% of the research content in 2023. Furthermore, all studies conducted in 2024 have utilized the decoder-only architecture. This trend indicates that in many current and future LLM4Security research and applications, the focus and resources will shift towards exploring and leveraging decoder-only architectures as the main approach. Looking at the total number of papers, researchers have begun to keenly focus on applying LLMs to solve challenging security tasks.
Applying LLMs to security. In our research, the use of LLMs can be categorized into agent-based processing and fine-tuning for specific tasks. Closed-source LLMs, represented by the GPT series, are the most popular in our studies. Researchers access LLMs online by calling APIs provided by LLM publishers and design task-specific prompts to guide LLMs to solve real-world problems with their training data (Jiang et al., 2020; Dong et al., 2023; Madaan et al., 2023), such as vulnerability repair and penetration testing (Deng et al., 2023a; Chen et al., 2023b; Wu et al., 2023a). Another approach involves locally fine-tuning open-source LLMs, by using datasets customized for specific functionalities, researchers are able to achieve significant performance improvements (Silva et al., 2024; Yang et al., 2023c).
In summary, the transition of LLMs in security, progressing from encoder-only architectures to decoder-only architectures, underscores the dynamic nature and flexibility of the field. This change has fundamentally altered the method for addressing security tasks, signaling ongoing innovation within the discipline.
RQ1 - Summary (1) We have gathered papers utilizing over 30 distinct LLMs for security tasks. These LLMs have been categorized into three groups based on their underlying architecture or principles: encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend in employing LLMs for security tasks, revealing that decoder-only architectures are the most prevalent. Specifically, over 15 LLMs fall into the decoder-only category, and a total of 92 papers have investigated the utilization of decoder-only LLMs in security tasks.
@@ 3.2. Trend Analysis
4. RQ2: What types of security tasks have been facilitated by LLM-based approaches?
@@ 4.1. Application of LLMs in Network Security
This section explores the application of LLMs in the field of network security. The tasks it includes web fuzzing, intrusion and anomaly detection, cyber threat analysis, and penetration testing.
@@ 4.2. Application of LLMs in Software and System Security
Bug repair. LLMs possess robust code generation capabilities, and their utilization in engineering for code generation can significantly enhance efficiency. However, code produced by LLMs often carries increased security risks, such as bugs and vulnerabilities (Perry et al., 2023). These program bugs can lead to persistent security vulnerabilities. Hence, automating the process of bug fixing is imperative, involving the use of automation technology to analyze flawed code and generate accurate patches to rectify identified issues. LLMs like CodeBERT (Le-Cong et al., 2023; Xia and Zhang, 2022; Huang et al., 2023; Zhang et al., 2023c), CodeT5 (Wang et al., 2023a; Tang et al., 2023; Huang et al., 2023), Codex (Jin et al., 2023; Fan et al., 2023; Xia and Zhang, 2023a), LLaMa (Tang et al., 2023), CodeLLaMa (Olausson et al., 2024; Silva et al., 2024) , CodeGEN (Xia and Zhang, 2023a), UniXcoder (Zhang et al., 2023c), T5 (Yuan et al., 2022), PLBART (Huang et al., 2023), and GPT Series (Zhang et al., 2023d; Olausson et al., 2024; Zhang et al., 2023c; Xia and Zhang, 2023b, a; Tang et al., 2023; Zhang et al., 2023h) have showcased effectiveness in generating syntactically accurate and contextually relevant code. This includes frameworks with encoder-decoder architecture like Repilot (Wei et al., 2023c), tailored specifically for producing repair patches. Utilizing LLMs for program repair can achieve competitive performance in producing patches for various types of errors and defects (Xia and Zhang, 2023b). These models effectively capture the underlying semantics and dependencies in code, resulting in precise and efficient patches. Moreover, fine-tuning LLMs on specific code repair datasets can further improve their ability to generate high-quality patches for real-world software projects. Integrating LLMs into program repair not only speeds up the error-fixing process but also allows software developers to focus on more complex tasks, thereby enhancing the reliability and maintainability of the software (Xia and Zhang, 2023a).
Several key findings summarize the research on LLMs for bug repair:
  • Interactive feedback: As demonstrated in the case of ChatGPT, notably enhances the accuracy of program repairs when integrated with interactive feedback loops (Xia and Zhang, 2023a). This iterative process of patch generation and validation fosters a nuanced comprehension of software semantics, thereby resulting in more impactful fixes.
  • Integration of external augmentation: By integrating domain-specific knowledge and technologies with the capabilities of LLMs, their performance is further enhanced. Custom prompts, fine-tuning for specific tasks, retrieving external data, and utilizing static analysis tools (Xia et al., 2023; Tang et al., 2023; Zhang et al., 2023b; Fu et al., 2022; Jin et al., 2023) significantly improve the effectiveness of bug fixes driven by LLMs.
@@ 4.5. Application of LLMs in Blockchain Security
RQ2 - Summary (1) We have divided cybersecurity tasks into six domains: software and system security, network security, information and content security, hardware security, and blockchain security. We have summarized the specific applications of LLMs in these domains. (2) We discussed 21 cybersecurity tasks and found that LLMs are most widely applied in the field of software and system security, with 76 papers covering 8 tasks. Only 5 papers mentioned the least applied domain—blockchain security.

### v3

@@ 4 The types and features of datasets used
3. RQ1: What types of security tasks have been facilitated by LLM-based approaches?
@@ 3.1. Application of LLMs in Network Security
This section explores the application of LLMs in the field of network security. The tasks include web fuzzing, intrusion and anomaly detection, cyber threat analysis, and penetration testing.
@@ 3.2. Application of LLMs in Software and System Security
Bug repair. LLMs possess robust code generation capabilities, and their utilization in engineering for code generation can significantly enhance efficiency. However, code produced by LLMs often carries increased security risks, such as bugs and vulnerabilities (Perry et al., 2023). These program bugs can lead to persistent security vulnerabilities. Hence, automating the process of bug fixing is imperative, involving the use of automation technology to analyze flawed code and generate accurate patches to rectify identified issues. LLMs like CodeBERT (Le-Cong et al., 2023; Xia and Zhang, 2022; Huang et al., 2023; Zhang et al., 2023c), CodeT5 (Wang et al., 2023a; Tang et al., 2023; Huang et al., 2023), Codex (Jin et al., 2023; Fan et al., 2023; Xia and Zhang, 2023a), LLaMa (Tang et al., 2023), CodeLLaMa (Olausson et al., 2024; Silva et al., 2024) , CodeGEN (Xia and Zhang, 2023a), UniXcoder (Zhang et al., 2023c), T5 (Yuan et al., 2022), PLBART (Huang et al., 2023), and GPT Series (Zhang et al., 2023d; Olausson et al., 2024; Zhang et al., 2023c; Xia and Zhang, 2023b, a; Tang et al., 2023; Zhang et al., 2023h) have showcased effectiveness in generating syntactically accurate and contextually relevant code. This includes frameworks with encoder-decoder architecture like Repilot (Wei et al., 2023c), tailored specifically for producing repair patches. Utilizing LLMs for program repair can achieve competitive performance in producing patches for various types of errors and defects (Xia and Zhang, 2023b). These models effectively capture the underlying semantics and dependencies in code, resulting in precise and efficient patches. Moreover, fine-tuning LLMs on specific code repair datasets can further improve their ability to generate high-quality patches for real-world software projects. Integrating LLMs into program repair not only speeds up the error-fixing process but also allows software developers to focus on more complex tasks, thereby enhancing the reliability and maintainability of the software (Xia and Zhang, 2023a). As demonstrated in the case of ChatGPT, notably enhances the accuracy of program repairs when integrated with interactive feedback loops (Xia and Zhang, 2023a). This iterative process of patch generation and validation fosters a nuanced comprehension of software semantics, thereby resulting in more impactful fixes. By integrating domain-specific knowledge and technologies with the capabilities of LLMs, their performance is further enhanced. Custom prompts, fine-tuning for specific tasks, retrieving external data, and utilizing static analysis tools (Xia et al., 2023; Tang et al., 2023; Zhang et al., 2023b; Fu et al., 2022; Jin et al., 2023) significantly improve the effectiveness of bug fixes driven by LLMs.
@@ 3.5. Application of LLMs in Blockchain Security
RQ1 - Summary (1) We have divided cybersecurity tasks into six domains: software and system security, network security, information and content security, hardware security, and blockchain security. We have summarized the specific applications of LLMs in these domains. (2) We discussed 21 cybersecurity tasks and found that LLMs are most widely applied in the field of software and system security, with 76 papers covering 8 tasks. Only 5 papers mentioned the least applied domain—blockchain security.
4. RQ2: What LLMs have been employed to support cybersecurity tasks?
4.1. Architecture of LLMs in Cybersecurity
Pre-trained Language Models (PLMs) have exhibited impressive capabilities across various NLP tasks (Shanahan, 2023; Kojima et al., 2023; Minaee et al., 2024; Wei et al., 2023a; Zhao et al., 2023). Researchers have noted substantial improvements in their performance as model size increases, with surpassing certain parameter thresholds leading to significant performance gains  (Shanahan, 2023; Hoffmann et al., 2022). The term ”Large Language Model” (LLM) distinguishes language models based on the size of their parameters, specifically referring to large-sized PLMs (Minaee et al., 2024; Zhao et al., 2023).However, there is no formal consensus in the academic community regarding the minimum parameter size for LLMs, as model capacity is intricately linked to training data size and overall computational resources (Kaplan et al., 2020). In this study, we adopt to the LLM categorization framework introduced by Panet et al. (Pan et al., 2024), which classifies the predominant LLMs explored in our research into three architectural categories: encoder-only, encoder-decoder, and decoder-only. We also considered whether the related models are open-source. Open-source models offer higher flexibility and can acquire new knowledge through fine-tuning on specific tasks based on pre-trained models, while closed-source models can be directly called via APIs, reducing hardware expenses. This taxonomy and relevant models are shown in Table 5. We analyzed the distribution of different LLM architectures applied in various cybersecurity domains, as shown in Fig 5.
Model Release Time Open Source
Encoder-Only BERT (8) 2018.10 Yes
RoBERTa (12) 2019.07 Yes
DistilBERT (3) 2019.10 Yes
CodeBERT (8) 2020.02 Yes
DeBERTa (1) 2020.06 Yes
GraphCodeBERT (1) 2020.09 Yes
CharBERT (1) 2020.11 Yes
Encoder-Decoder T5 (4) 2019.10 Yes
BART (1) 2019.10 Yes
PLBART (3) 2021.03 Yes
CodeT5 (5) 2021.09 Yes
UniXcoder (1) 2022.03 Yes
Flan-T5 (1) 2022.10 Yes
Decoder-Only GPT-2 (9) 2019.02 Yes
GPT-3 (4) 2020.04 Yes
GPT-Neo (1) 2021.03 Yes
CodeX (9) 2021.07 No
CodeGen (5) 2022.03 Yes
InCoder (1) 2022.04 Yes
PaLM (3) 2022.04 No
Jurassic-1 (1) 2022.04 No
GPT-3.5 (52) 2022.11 No
LLaMa (4) 2023.02 Yes
GPT-4 (38) 2023.03 No
Bard (8) 2023.03 No
Claude (3) 2023.03 No
StarCoder (3) 2023.05 Yes
Falcon (2) 2023.06 Yes
CodeLLaMa (4) 2023.08 Yes
Encoder-only LLMs. Encoder-only models, as their name implies, comprise solely an encoder network. Initially designed for language understanding tasks like text classification, these models, such as BERT and its variants (Aghaei et al., 2023b; He et al., 2021; Devlin et al., 2019; Liu et al., 2019; Ma et al., 2020; Feng et al., 2020; Guo et al., 2021; Sanh et al., 2020), aim to predict a class label for input text (Devlin et al., 2019). For instance, BERT, which adopts the encoder architecture of the Transformer model, is mentioned in 35 papers included in this study. Encoder-only LLMs use a bidirectional multi-layer self-attention mechanism to calculate the relevance of each token with all other tokens, thereby capturing semantic features that include the global context. This architecture is mainly used for processing input data, focusing on understanding and encoding information rather than generating new text. Researchers employed these models to generate embeddings for data that is relavent to cybersecurity (such as traffic data and code), mapping complex data types into vector space. These models typically use a masking strategy during pre-training, and the complex training strategies increase training time and the risk of overfitting. In the realm of cybersecurity, researchers have adopted advanced models that offer capabilities much needed in cybersecurity tasks such as code understanding (Wang et al., 2023c) and traffic analysis (Aghaei and Al-Shaer, 2023).
Various prominent models, including CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), RoBERTa (Liu et al., 2019), CharBERT (Ma et al., 2020), DeBERTa (He et al., 2021), and DistilBERT (Sanh et al., 2020), have gained widespread usage due to their ability to effectively process and analyze code, making them valuable tools in the field of cybersecurity. An example is RoBERTa (Liu et al., 2019), which enhances BERT’s robustness through various model design adjustments and training techniques. These include altering key hyperparameters, eliminating the next-sentence pre-training objective, and utilizing substantially larger mini-batches and learning rates during training. CodeBERT (Feng et al., 2020) is a bimodal extension of BERT that utilizes both natural language and source code as its input. It employs a replaced token detection task to bolster its understanding of programming languages, in order to tackle code generation and vulnerability detection tasks. The encoder-only architecture provides models with excellent data representation capabilities. Note that these aforementioned BERT variants were not initially designed for cybersecurity tasks. Instead, their application in the cybersecurity field stems from their capabilities as general models in NLP tasks for code semantics interpretation and understanding. In contrast, SecureBERT (Aghaei et al., 2023b) is a BERT variant specifically designed for cyber threat analysis tasks. Its development highlights the robustness and flexibility of encoder-only architecture models across different tasks. Diverse training tasks and specialized training schemes enhance the model’s feature representation capabilities and boosts its performance in cybersecurity-related tasks.
Regarding the model applicability, as shown in the Figure 5, encoder-only models initially garnered attention in the fields of network cybersecurity (Alkhatib et al., 2022) and software and systems cybersecurity (Lee et al., 2023; Xia and Zhang, 2022). In 2023, this concept was extended to the field of information and content cybersecurity, utilizing encoder-only models to harmful content on social media platforms (Mets et al., 2023; Hanley and Durumeric, 2023; Cai et al., 2024).
Figure 5. Distribution and trend of different model architectures.
Encoder-decoder LLMs. The Transformer model, based on the encoder-decoder architecture (Vaswani et al., 2023), consists of two sets of Transformer blocks: the encoder and decoder. Stacked multi-head self-attention layers are used by the encoder to encode the input sequence, generating latent representations. In contrast, the decoder performs cross-attention on these representations and sequentially produces the target sequence. The structure of encoder-decoder LLMs makes them highly suitable for sequence-to-sequence tasks such as code translation and summarization. However, their complex architecture requires more computational resources and high-quality labeled data.
Models like BART (Lewis et al., 2019), T5 (Raffel et al., 2023), and CodeT5 (Wang et al., 2021) exemplify this architecture. CodeT5 (Wang et al., 2021) and PLBART (Ahmad et al., 2021) have built upon the foundation of their original models by introducing bimodal inputs of programming language and text, demonstrating effective code comprehension capabilities. Raffle et al. (Raffel et al., 2023) show that almost all NLP tasks can be framed as a sequence-to-sequence generation task in their work. In LLM4Security, the encoder-decoder architecture was first attempted to be applied in the field of network security (Liu et al., 2020a). However, subsequent research has not widely adopted this approach, possibly due to the complexity of the encoder-decoder structure. From another perspective, owing to its flexible training strategy and excellent adaptability to complex tasks, the encoder-decoder model was later extended to other cybersecurity tasks such as program fuzzing (Deng et al., 2023c), reverse engineering (Armengol-Estapé et al., 2024), and phishing emails detection (Jamal and Wimmer, 2023).
Decoder-only LLMs. Unlike the encoder-decoder architecture, which involves the encoder processing input text and the decoder generating output text by predicting subsequent tokens from an initial state, decoder-only LLMs rely solely on the decoder module to produce the target output text (Radford et al., 2018). This autoregressive training paradigm allows decoder-only models to generate longer-form outputs token by token, making them well-suited for producing detailed analyses, advisories, and even code relevant to cybersecurity. The attention mechanism in these models also enables them to flexibly draw upon the extensive knowledge stored in their parameters and apply it to the current context.
GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), GPT-3.5 (OpenAI, 2022), and GPT-4 (OpenAI, 2023b) belong to the GPT series of models, among which GPT-3.5 and GPT-4 are the models most frequently used to address various cybersecurity issues in this study, covering almost all cybersecurity applications (Ferrag et al., 2024; Fayyazi and Yang, 2023; Scanlon et al., 2023; Deng et al., 2023a). Their strong few-shot learning abilities allow rapid development of new cybersecurity capabilities with minimal fine-tuning. More specialized versions like Codex (Chen et al., 2021) and others have been fine-tuned for specific code-related tasks. Open-source models like GPT-Neo (Black et al., 2021), LLaMa (Touvron et al., 2023), and Falcon (Penedo et al., 2023) also follow this architecture. Additionally, code generation LLMs such as CodeGen (Nijkamp et al., 2023), InCoder (Fried et al., 2023), StarCoder (Li et al., 2023a), and CodeLLaMa (Rozière et al., 2024) have been widely used for bug detection and repair, as well as for vulnerability repair (Wu et al., 2023a; Xia and Zhang, 2023b; Yang et al., 2023c). The large context window of decoder-only models allows them to take in and utilize more context about the cybersecurity task, like related vulnerabilities, reports, and code snippets.
Due to the powerful natural language generation capabilities of the decoder-only architecture, researchers initially attempted to apply it to the generation of fake cyber threat intelligence (Ranade et al., 2021). Decoder-only LLMs have gained prominence in recent years, especially in 2022 and 2023 as shown in Figure 5, witnessing a surge in development and commercial adoption by leading Internet companies. For instance, Google introduced Bard (Google, 2023), while Meta unveiled LLaMa (Touvron et al., 2023). Unlike GPT-4 and its derivative application, ChatGPT, which quickly found integration into various cybersecurity tasks. These newer models have yet to see widespread adoption in the cybersecurity domain.
4.2. Trend Analysis
Illustrated in Figure 5, from 2020 to 2024, there have been significant shifts in the preference and utilization of LLM architectures across cybersecurity tasks. The selection of decoder-only, encoder-decoder, and encoder-only structures has influenced diverse research directions and solutions in the cybersecurity field. This examination delves into the trends regarding the adoption of these architectures over time, reflecting the evolving landscape of LLM applications for cybersecurity tasks.
Table 6. Overview of the distribution of LLMs in the open-source community.
(a) Top 20 most downloaded models on Huggingface.
Model Architecture
BERT-base Encoder-only
DistilBERT-base Encoder-only
GPT2 Decoder-only
RoBERTa-large Encoder-only
RoBERTa-base Encoder-only
xlm-RoBERTa-large Encoder-only
xlm-RoBERTa-base Encoder-only
DeBERTa-base Encoder-only
Qwen-VL-Chat Decoder-only
T5-small Decoder-encoder
BERT-base-cased Encoder-only
T5-base Decoder-encoder
BERT-base-uncased Encoder-only
CamemBERT-base Encoder-only
DistilGPT2 Decoder-only
DistilRoBERTa-base Encoder-only
LLaMa3-8B Decoder-only
ALBERT-base-v2 Encoder-only
DeBERTa-v3-base Encoder-only
ByT5-small Decoder-encoder
(b) Top 20 most liked models on Huggingface.
Model Architecture
BLOOM-176B Decoder-only
LLaMa3-8B Decoder-only
LLaMa2-7B Decoder-only
Mixtral-8x7B Decoder-only
Mixtral-7B Decoder-only
Phi-2 Decoder-encoder
Gemma-7B Decoder-only
ChatGLM-6B Decoder-only
StarCoder Decoder-only
Falcon-40B Decoder-only
Grok-1 Decoder-only
ChatGLM2-6B Decoder-only
GPT2 Decoder-only
Dolly-v2-12B Decoder-only
BERT-base Encoder-only
Zephyr-7B Decoder-only
OpenELM Decoder-only
Phi-1.5 Decoder-encoder
Yi-34B Decoder-only
Flan-T5 Decoder-encoder
Timeline and Model Architecture distribution. In 2020 and 2021, the use of LLMs in cybersecurity was limited, with only 3 research papers exploring their potential. In 2020, encoder-decoder LLMs, known for their strong performance on sequence-to-sequence tasks, were the sole architecture used in a single paper. However, in 2021, the focus shifted to decoder-only LLMs, which excel at generating longer-form outputs and handling diverse queries due to their autoregressive generation capabilities and large context windows. This shift can be attributed to the research emphasis on LLM performance in natural language processing tasks and innovations in LLM architectures during this period (Kaplan et al., 2020; Brown et al., 2020).
The year 2022 marked a significant turning point, with the number of papers employing LLMs for cybersecurity tasks surging to 11, surpassing the combined total from the previous two years. This year also saw increased diversity in the LLM architectures used. Encoder-only LLMs, valued for their representation learning and classification abilities, were utilized in 46% of the research (5 papers). Encoder-decoder LLMs, with their strong performance on well-defined tasks, were featured in 18% (2 papers), while decoder-only LLMs, leveraging their knowledge recall and few-shot learning capabilities, garnered 36% of the research interest (4 papers). This varied distribution highlights the active exploration of different architectures to address the diverse needs and challenges in cybersecurity.
The years 2023 and 2024 witnessed a significant shift towards decoder-only LLMs, which emerged as the primary architecture for addressing cybersecurity challenges. This trend is closely tied to the powerful text comprehension, reasoning capabilities (Ouyang et al., 2022; Wei et al., 2023b), and open-ended generation demonstrated by chatbots like ChatGPT. These decoder-only models require minimal fine-tuning and can generate both syntactically correct and functionally relevant code snippets (Laskar et al., 2023a; Sadik et al., 2023). In 2023, decoder-only LLMs accounted for 68.9% of the total research, while encoder-decoder LLMs and encoder-only LLMs contributed 10.7% (14 papers) and 22.1% (27 papers), respectively. Remarkably, all studies conducted in 2024 utilized the decoder-only architecture, indicating a strong focus on exploring and leveraging the unique advantages of these models in cybersecurity research and applications.
The dominance of decoder-only LLMs in cybersecurity research aligns with the broader trends in the LLM community. An analysis of the top 20 most liked and downloaded LLMs on Huggingface (hug, 2023), a popular open-source model community, reveals that while encoder-only models like BERT and its variants have the highest number of downloads, decoder-only models are gaining significant traction. Moreover, 16 out of the top 20 most liked LLMs are decoder-only models, indicating a strong preference and excitement for their potential to handle complex, open-ended tasks. The growing interest in decoder-only LLMs can be attributed to their strong generation, knowledge, and few-shot learning abilities, which make them well-suited for the diverse challenges in cybersecurity. However, the larger parameter size of these models compared to encoder-only models may limit their current adoption due to the scarcity of computational resources (Fedus et al., 2022).
Applying LLMs to cybersecurity. In our research, the use of LLMs can be categorized into agent-based processing and fine-tuning for specific tasks. Closed-source LLMs, represented by the GPT series, are the most popular in our studies. Researchers access LLMs online by calling APIs provided by LLM publishers and design task-specific prompts to guide LLMs to solve real-world problems with their training data (Jiang et al., 2020; Dong et al., 2023; Madaan et al., 2023), such as vulnerability repair and penetration testing (Deng et al., 2023a; Chen et al., 2023b; Wu et al., 2023a). Another approach involves locally fine-tuning open-source LLMs, by using datasets customized for specific functionalities, where researchers are able to achieve significant performance improvements (Silva et al., 2024; Yang et al., 2023c).
In summary, the transition of LLMs in cybersecurity, progressing from encoder-only architectures to decoder-only architectures, underscores the dynamic nature and flexibility of the field. This change has fundamentally altered the method for addressing cybersecurity tasks, signaling ongoing innovation within the discipline.
RQ2 - Summary (1) We have gathered papers utilizing over 30 distinct LLMs for cybersecurity tasks. These LLMs have been categorized into three groups based on their underlying architecture or principles: encoder-only, encoder-decoder, and decoder-only LLMs. (2) We analyzed the trend in employing LLMs for cybersecurity tasks, revealing that decoder-only architectures are the most prevalent. Specifically, over 15 LLMs fall into the decoder-only category, and a total of 98 papers have investigated the utilization of decoder-only LLMs in cybersecurity tasks.